the base of a triangle is 6 meters less than the height. the area is 12 meters squared. find the base and the height of the triangle.
Answer by stanbon(75887)

the base of a triangle is 6 meters less than the height. the area is 12 meters squared. find the base and the height of the triangle.
---------------------------------------------------------
Area = (1/2)base*height
12 = (1/2)(h-6)h
24 = h^2-6h
---
Rearrange:
h^2-6h-24 = 0
h = [6+- sqrt(36-4*-24)]/2
---
h = [6+-sqrt(132)]/2
Positive solution:
h = [6 + 2sqrt(33)]/2
height = 3+sqrt(33) meters
---
base = height-6 = -3+sqrt(33) meters
===========================================
Cheers,
Stan H.